OpenStudy (anonymous):

Solve for y: 3x + 2y < 12. Answer y < -3/2x + 6 y < -2/3x + 4 y > 3/2 x + 6 y < 3/2x + 4

OpenStudy (anonymous):

3x+2y=12 2y=-3x+12 y=-3x+12/2 y=-3x/2+6

OpenStudy (ryan123345):